OpenStudy (anonymous):

Given a wavelength of 121,270 millimeters, what is the wavelength in kilometers? 0.121270 12.1270 1,212.70 12,127.0

jimthompson5910 (jim_thompson5910):

hint: 1000 millimeters = 1 meter 1000*1000 millimeters = 1000*1 meter ... Multiply both sides by 1000. 1,000,000 millimeters = 1000 meters 1,000,000 millimeters = 1 kilometer

OpenStudy (whpalmer4):

121270 mm * (1 km / 1000 m) * (1 m / 1000 mm) = _______ km write it with ratios like that and you'll catch any errors you make where you don't apply enough conversions, or apply them in reverse...just make sure the units cancel properly.